Simplify 2|8|+3|3-12|+1/|12|-4|2(5-1)| Thanks for your help.
Answer by macston(5194)

.
2|8|+3|3-12|+1/|12|-4|2(5-1)|
2(8)+3(9)+(1/12)-(4*2*4)=16+27+1/12-32=133/12
ANSWER: 133/12